Reports 1-1 of 1 Clear search Modify search
AdV-INJ (Input Power Control (IPC) system)
derossi - 15:08 Friday 09 August 2019 (46631) Print this report
laser power to the ITF decreased

Upon request of ISC I decreased the laser power (with EIB IPC 1) from 20 W to 19.4 W.

 

Images attached to this report
Comments to this report:
mantovani, ruggi, tacca - 11:51 Sunday 11 August 2019 (46646) Print this report
We spent Friday morning to try to restore a stable configuration of the ITF without success: it was impossible to have a good CMRF and close the AA BS loop at the same time. Therefore, we decided to check the input power during the last fine tuning of the CO2 lasers (Jun 28th, entry 46226) and we discovered that it was around 19.5 W. Going back to this value helped a lot and it was quite easy to recover a stable interferometer with good CMRF and all the loops properly closed.
Search Help
×

Warning

×